### v3

# GenMix: Effective Data Augmentation with Generative Diffusion Model Image Editing

Abstract
Data augmentation is widely used to enhance generalization in visual classification tasks. However, traditional methods struggle when source and target domains differ, as in domain adaptation, due to their inability to address domain gaps. This paper introduces GenMix, a generalizable prompt-guided generative data augmentation approach that enhances both in-domain and cross-domain image classification. Our technique leverages image editing to generate augmented images based on custom conditional prompts, designed specifically for each problem type. By blending portions of the input image with its edited generative counterpart and incorporating fractal patterns, our approach mitigates unrealistic images and label ambiguity, improving performance and adversarial robustness of the resulting models. Efficacy of our method is established with extensive experiments on eight public datasets for general and fine-grained classification, in both in-domain and cross-domain settings. Additionally, we demonstrate performance improvements for self-supervised learning, learning with data scarcity, and adversarial robustness. As compared to the existing state-of-the-art methods, our technique achieves stronger performance across the board.
Index Terms:
Data Augmentation, Mixup Data Augmentation, Diffusion Models, Generative Synthetic Data, Synthetic Data
1 Introduction
Deep Neural Networks (DNNs) have achieved remarkable advancements across various computer vision tasks, such as image classification Dosovitskiy et al. (2020); He et al. (2016a); Liu et al. (2022); Tan and Le (2019), object detection Fang et al. (2021); He et al. (2017); Wang et al. (2021), image captioning Fei et al. (2022); Ke et al. (2019), human pose estimation Liu et al. (2021); Luo et al. (2021), and image segmentation Badrinarayanan et al. (2017); Strudel et al. (2021). However, as the complexity of these tasks increases, so do the computational costs associated with modern DNN architectures, including larger memory footprints, model sizes, higher FLOPs, and longer inference times DeVries and Taylor (2017). Recent models employ billions of parameters to enhance feature representation. This often leads to overfitting and challenges in generalization, especially when the training data is limited or poorly structured. These issues become even more pronounced in cross-domain settings, where models must adapt to divergent data distributions.
Image mixing based data augmentation techniques have proven effective for enhancing the generalization of deep learning models. A wide array of such techniques has been introduced, including CutMix Yun et al. (2019a), Mixup Kim et al. (2020a), SnapMix Huang et al. (2021), Co-Mixup Kim et al. (2020b), SaliencyMix Uddin et al. (2020), GuidedMixup Kang and Kim (2023), MixPro Zhao et al. (2022), TransMix Chen et al. (2022a), and PuzzleMix Kim et al. (2020a). These methods typically involve combining portions of randomly selected images of corresponding labels through various mixing strategies to generate augmented samples - see top row in Fig. 1. By employing linear interpolation or other blending techniques, they create novel training images to enrich the dataset for addressing potential model overfitting Yun et al. (2019b); Zhang et al. (2018); Han et al. (2022); Qin et al. (2023); Huang et al. (2021); Islam (2022).
Though effective, image mixing based data augmentation encounters challenges, such as the omission of salient regions and label ambiguities due to the random blending of image content Kim et al. (2020a). To mitigate these issues, several approaches have incorporated saliency-driven methods, which prioritize the preservation of important regions by overlaying them onto less significant background areas of the target image Kim et al. (2020a); Uddin et al. (2020); Kang and Kim (2023); Islam et al. (2022). While these techniques offer improvements, they introduce considerable computational overhead and rely on the accuracy of the saliency methods, which is a limiting factor. Furthermore, by combining images from diverse categories, these techniques can still overlook crucial contextual information, leaving the problem of inadequate feature preservation unresolved.
Figure 1: Representative examples of augmentation techniques. Top: Conventional image mixing methods like CutMix Yun et al. (2019b), Mixup Zhang et al. (2018), GridMix Baek et al. (2021), SmoothMix Lee et al. (2020), ResizeMix Qin et al. (2020) and FMix Harris et al. (2020) use interpolation or other strategies to mix source and target images, largely disregarding the natural appearance of the resulting image. Bottom: Our GMix is a label-preserving technique like AugMix Hendrycks et al. (2019), PixMix Hendrycks et al. (2022) and IPMix Huang et al. (2023). However, it utilizes a generative diffusion model with custom tailored prompts for high quality augmentation. Three different outputs (1A, 2A and 3A) for the target are shown for GenMix.
Diffusion Models (DMs) Takagi and Nishimoto (2023); Du et al. (2023); Luo et al. (2023); Saharia et al. (2022); Dhariwal and Nichol (2021) have recently emerged as powerful tools for image-to-image generation and editing. Some studies Trabucco et al. (2024); Azizi et al. (2023) have also explored using DM generated images to augment training data. However, we empirically observe limited performance gains from these approaches, with models trained on synthetic data occasionally underperforming compared to the baseline models without augmentation Azizi et al. (2023). This can be attributed to the limited control over image generation of the DMs. Due to the model sensitivity to conditional prompts, generating complex scenes, layouts, and shapes becomes challenging Zhang et al. (2023). Poorly crafted prompts often yield unsuitable images for augmentation, which can lead to model adaption to an incorrect data distribution. This underscores the need for more precise prompt selection and better control mechanisms in the augmentation process to improve both in-domain generalization and cross-domain performance.
To address this, we propose GenMix, a generlizable data augmentation technique that leverages generative pre-trained diffusion models to compute diverse augmentation samples through carefully designed conditional prompts. GMix creates hybrid images by blending original and generated content, maintaining essential semantics of the original image, see bottom row of Fig. 1. To further enrich structural diversity, we incorporate self-similarity fractals into the hybrid images, a strategy proven to enhance machine learning safety Hendrycks et al. (2022); Huang et al. (2024). This blending mitigates overfitting and improves model performance. Our strategy is unlike any previous attempts of directly using DM images for augmentation Azizi et al. (2023); Trabucco et al. (2024). It effectively incorporates a range of operations in its components to achieve outstanding eventual performance - see Table I. Our experiments show that GMix surpasses current state-of-the-art (SOTA) augmentation methods by improving generalization and adversarial robustness. Additionally, it facilitates prompt-guided learning in both in-domain and cross-domain tasks, making it highly adaptable for diverse computer vision problems. Furthermore, our results confirm that GMix is compatible with a wide range of datasets and it integrates seamlessly into existing SOTA frameworks.
The main contributions of this work are as follows:
  • We introduce a novel data augmentation method, GenMix that leverages image-editing with generative diffusion models guided by curated conditional prompts. Our technique merges original and generated images, followed by blending self-similarity fractals, which reduces overfitting and ensures robust training.
  • We explore the applicability of GMix to both in-domain classification and domain adaptation tasks, introducing sets of bespoke conditional prompt for enhanced adaptability. Notably, GMix is the first method to address both in-domain and domain adaptation tasks effectively.
  • We demonstrate with extensive experiments across eight datasets that GMix outperforms existing SOTA augmentation methods in general classification, fine-grained classification, adversarial robustness, and cross-domain tasks.
TABLE I: Comparison of data augmentation methods. Components include central operations and inputs. Tasks are the key problems for which the efficacy of the methods is established.
Input
CutMix
Mixup
GridMix
SMix
ResMix
FMix
AugMix
PixMix
IPMix
Diff-Mix
Aziz
DA-Fusion
Ours
Components Source image
Target image
Fractal image
Textual prompts
Specific prompts
Interpolation
Concatenation
Tasks Adversarial robustness
General classification
Fine-grained
Transfer learning
Data scarcity
Open-partial
Open-set
Partial-set
A preliminary version of this research was presented at IEEE CVPR 2024 Islam et al. (2024). This article is a major extension of the preliminary work. Here, we propose a more effective approach to fuse the original and generated images by introducing a smoothness factor, leading to consistent performance gains. As compared to Islam et al. (2024), we also extend the fractal image dataset to add more variety and structural complexity to the the final augmented images. Moreover, besides providing considerable further details for the technique, prior art and evaluation, we also explore the applicability of our work to domain adaption tasks such as open-partial, open-set, and partial-set to facilitate a more comprehensive view of the resulting data augmentation. To the best of our knowledge, GenMix is the first augmentation technique that enables significant performance gains in both in-domain and cross-domain tasks, signifying its importance in training robust large-scale learning systems.
2 Related Work
Data augmentation is widely adopted to enhance the diversity of training datasets. Traditional methods apply basic transformations, such as flipping, rotation, and color jittering, to create new images from the original data. However, while these transformations can increase performance and improve generalization, they often lack semantic diversity, failing to capture key features critical for more complex tasks. This motivates the need for more advanced augmentation techniques that can introduce greater variation in the data while preserving essential content.
Diffusion Models for Augmentation. Recently, there has been a growing interest in leveraging Diffusion Models (DMs) for data augmentation. Azizi et al. Azizi et al. (2023) proposed using text-to-image fine-tuned diffusion models to generate synthetic samples for ImageNet classification, demonstrating that incorporating these generated images into the training set can enhance classification performance. Similarly, Trabucco et al. Trabucco et al. (2024) utilized off-the-shelf diffusion models to create diverse and semantically rich images via prompt engineering, aiming to improve image classification on various datasets. Furthermore, Li et al. Li et al. (2023) explored using diffusion models for data augmentation in conjunction with knowledge distillation, particularly in scenarios where real images are unavailable. These efforts underscore the potential of diffusion models to enrich data augmentation techniques.
Image Mixing for Augmentation. Image mixing has become a widely employed category of data augmentation techniques, enhancing the performance and robustness of deep learning models Liang et al. (2023); Yan et al. (2023); Mensink and Mettes (2023); Chen and Lu (2023). Notable state-of-the-art methods in this category include CutMix, AugMix and PuzzleMix. CutMix Yun et al. (2019a) improves out-of-distribution generalization by randomly overlaying a patch from a source image onto a target image. AugMix Hendrycks et al. (2019) stochastically combines augmentation operations like equalize, posterize, and translate-x to introduce semantic diversity. PuzzleMix Kim et al. (2020a) refines mixup techniques by considering image saliency and local statistics. Similarly, another technique SaliencyMix Uddin et al. (2020) leverages saliency maps to focus on the most important image regions, preserving their integrity. Manifold Mixup Verma et al. (2019) blends hidden network states during training, creating an interpolated state for improved feature representation. Another well-known method is PixMix Hendrycks et al. (2022) that extends AugMix Hendrycks et al. (2019) by mixing input images with fractal and feature visualization images, improving machine learning safety. A comprehensive overview of various image-mixing methods, along with their components and task accomplishment is presented in Table I.
Automated Data Augmentation. Several researchers have also explored automated data augmentation for improving model performance. For instance, AutoAugment Cubuk et al. (2018) utilizes reinforcement learning to discover optimal data augmentation strategies, while RandAugment Cubuk et al. (2020) employs a set of randomized augmentation operations to enhance model generalization and robustness. AdaAug Cheung and Yeung (2021) has been introduced to efficiently learn adaptive augmentation policies that vary depending on the class or even the instance.
Although existing methods are effective within their settings, they often lack in generalizability. The conventional image mixing techniques remain brittle due to their non-adaptive nature while more recent generative modeling based methods underperform due to poorly crafted prompts. Moreover, the desired augmentation image formation by these methods largely ignore the semantics while blending the available images. Unlike these previous techniques, our method focuses on combining original and generated images in a more sophisticated manner for semantic preservation while also leveraging a predefined library of conditional prompts. Our hybrid images are also blended with fractal patterns to further boost the overall model generalization and robustness.
Figure 2: Overview of the proposed GMix method. An input image and a randomly selected prompt are input to a pre-trained image editing diffusion model to obtain a generated image. Input and generated images are smoothly concatenated using a binary mask to obtain a hybrid image. A random fractal image is finally blended with this hybrid image to obtain the augmented image.
3 Proposed Method
3.1 Overview
Existing image mixing based data augmentation techniques may induce label ambiguities by pasting the source image contents onto the target image and consequently replacing either important regions of the foreground or the required background context Yun et al. (2019b); Kang and Kim (2023). The central idea of our approach is to augment the source image only while preserving the key semantics and concatenating a portion of the source image with its generative counterpart. At the same time, our method introduces diverse object details and contexts in the resulting image for effective augmentation.
The schematics of the proposed GMix technique is provided in Fig. 2. Our method comprises three main steps: prompt-guided image editing, seamless concatenation, and fractal interpolation. Initially, we use conditional prompts with a diffusion-based image editing module to compute generative counterparts of the source image. Then, a portion of the source image is concatenated with the rest of the generated image, forming a hybrid image. This approach ensures that the training model has access to the original visual content along with the generated one. Finally, the hybrid image is blended with a random fractal image, obtaining the final training image with a diverse structure. Improvements of our image by blending fractal is inspired by Hendrycks et al. (2022, 2021). However, in our case, the blending objective for augmentation is unique to our approach.
3.2 GenMix Details
The proposed GMix is quite effective data augmentation method that aims to enhance the generalization and robustness of the deep neural networks. Formally, $I_{i}\in\mathbb{R}^{h\times w\times c}$ is a training dataset image. Our data augmentation method is defined as: $\mathcal{D}$mix$(\cdot):\mathbb{R}^{h\times w\times c}\to\mathbb{R}^{h\times w\times c}$. Input image $I_{i}$ goes through proposed generation using randomly selected prompt $p_{j}$, seamless concatenation using random mask $M_{u}$, and linearly interpolated using a random fractal image $F_{v}$ to obtain the final augmented image $A_{ijuv}$. The overall augmentation process, as shown in Algorithm 1, may be represented as
$A_{ijuv}=\mathcal{D}_{\text{mix}}(I_{i},p_{j},M_{u},F_{v},\lambda).$ (1)
Prompt Guided Image Editing: Our image editing step $\mathcal{E}(.)$ consists of a pre-trained image editing diffusion model that requires a prompt $p_{j}$ from a predefined set of $k$ prompts, $P=\{p_{1},p_{2},\dots,p_{k}\}$ where $j\in[1,k]$ along with the input training image $I_{i}$ and produces an edited counterpart image $\hat{I}_{ij}$. The image generation process in conventional diffusion models is often open-ended for image-to-image or text-to-image translation. On the other hand, image editing is guided by textual prompts. The aim is to obtain diverse image-to-image translations guided by textual prompts. In our case, as the goal is to achieve a slightly modified but not too different version of $I_{i}$, filter-like prompts are curated in $\mathcal{P}$ which do not alter the image drastically. Examples of the prompts used in GMix are shown in Figure 3. The overall generation step can be represented as:
$\hat{I}_{ij}=\mathcal{G}(I_{i},p_{j}),\text{ where }p_{j}\in\mathcal{P}$ (2)
Seamless Concatenation: We concatenate a portion of the original input image $I_{i}$ with its edited version $\hat{I}_{ij}$ using a randomly selected mask $M_{u}$ from the set of masks to create a hybrid image $H_{iju}$:
$H_{iju}=(\hat{I}_{ij}\odot M_{u})+(I_{i}\odot(\mathbb{1}-M_{u})),$ (3)
where $M_{u}$ is a mask, $\odot$ is a pixel-wise multiplication operator. The set of masks contains four kinds of masks including horizontal, vertical and flipped versions. Such masking ensures the availability of the semantics of the input image to the learning network while reaping the benefits of the generated images.
The mask $M_{u}$ consists of zeros, ones, and smoothly varying values for seamless concatenation. To create a smooth transition between two images, our mask consists of three parts: $M_{0}\in\mathbb{R}^{h\times(w-b)/2}$ containing zeros, $M_{1}\in\mathbb{R}^{h\times(w-b)/2}$ containing ones, and $M_{b}\in\mathbb{R}^{h\times b}$ containing the blending matrix where $b$ is the blending width. These three parts are stacked to form the mask: $M_{u}=\begin{bmatrix}M_{0}&M_{b}&M_{1}\end{bmatrix}$ either horizontally or vertically. We generate the blending matrix $M_{b}$ such that the values smoothly transition from 0 on the $M_{0}$ side to 1 on the $M_{1}$ side, over the specified blending width $b$. The same $M_{u}$ is applied across all three color channels of one image and (1-$M_{u}$) is applied to the second image and then both images are added to get the hybrid image as shown by Eq. (3).
Fractal Interpolation: A self-similarity fractals dataset $\mathcal{F}$ is collected and used for inducing structural variations in the hybrid images. The interpolated image $H_{iju}$ is linearly computed between a randomly selected fractal image $F_{v}\in\mathcal{F}$ and the hybrid image $H_{iju}$:
$A_{ijuv}=\lambda F_{v}+(1-\lambda)H_{iju},\text{ }0\leq\lambda<1$ (4)
where $\lambda$ is the interpolation factor. This results in the final augmented image $A_{ijuv}$ used to train or fine-tune deep neural networks. The overall augmentation process of GMix can be represented as:
$A_{ijuv}=(1-\lambda)(I_{i}\odot M_{u}+\hat{I}_{ij}\odot(\mathbb{1}-M_{u}))+\lambda F_{v},$ (5)
An ablation study is conducted to find the appropriate value of $\lambda$.
Algorithm 1 GMix
1:$I_{i}\in\mathcal{D}$ training images dataset, $m$: number of augmented images, $p_{j}\in\mathcal{P}$ set of textual prompts, $M_{u}\in\mathcal{M}$ set of smooth masks, $F_{v}\in\mathcal{F}$ set of self-similarity fractal images, $\lambda$: interpolation weight
2:$\mathcal{D}^{\prime}$: $m$ Augmented images
3:$\mathcal{D}^{\prime}\leftarrow\emptyset$
4:for each image $I_{i}$ in $\mathcal{D}$ do
5:     for $a$ in $\{1:m\}$ do
6:         Randomly select prompt $p_{j}$ from $\mathcal{P}$
7:         Image Editing: $\hat{I}_{ij}\leftarrow\mathcal{E}(I_{i},p_{j})$
8:         Randomly select mask $M_{u}$ from $\mathcal{M}$
9:         Concatenation: $H_{iju}\leftarrow$ $M_{u}\odot I_{i}+(1-M_{u})\odot\hat{I}_{ij}$
10:         Randomly select $F_{v}$ from $\mathcal{F}$
11:         Linear Interpolation: $A_{ijuv}\leftarrow(1-\lambda)H_{iju}+\lambda F_{v}$
12:         Add $A_{ijuv}$ to $\mathcal{D}^{\prime}$
13:     end for
14:end for
15:return $D^{\prime}$
Figure 3: Different bespoke conditional prompts - columns - are used to obtain generated images preserving important features and adding rich visual appearance to the input images. Each row shows a representative input from a different dataset.
4 Experimental Setup
Datasets. To enable comparisons with previous research on image augmentation Yun et al. (2019b); Zhang et al. (2018); Kang and Kim (2023); Uddin et al. (2020); Hendrycks et al. (2019); Cheung and Yeung (2021); Cubuk et al. (2020, 2019); Kim et al. (2020a, b); Huang et al. (2021); Lim et al. (2019); Verma et al. (2019), we assess our approach using both general image classification and fine-grained image classification datasets. For general image classification, we use three datasets: ImageNet-1K Deng et al. (2009), CIFAR100 Krizhevsky et al. (2009), and Tiny-ImageNet-200 Le and Yang (2015). For fine-grained image classification, we employ three datasets: CUB Nilsback and Zisserman (2008), Stanford Cars Krause et al. (2013), and Aircraft Maji et al. (2013). Moreover, we use Flower102 Nilsback and Zisserman (2008), Aircraft Maji et al. (2013), and Stanford Cars Krause et al. (2013) for evaluation related to Transfer Learning. The data scarcity related experiments are conducted with Flower102 Nilsback and Zisserman (2008) dataset. Our experiments for domain adaption additionally use OfficeHome Venkateswara et al. (2017) dataset. These datasets encompass a variety of scenarios, featuring images with a broad spectrum of objects like plants, animals, vehicles, human actions, and general objects across diverse scenes and textures.
Implementation Details. For image editing, we employed InstructPix2Pix pre-trained image editing diffusion model Brooks et al. (2023) utilizing our custom-built textual prompts. To create the Mask $M_{u}$ we used a fixed blending width of 20 pixels. We use $\lambda=0.20$ for blending the fractal image in Eq. (4).
Textual Prompt Selection. To ensure that only suitable prompts are used for classification, a custom textual library of filter-like global visual effects is predefined, comprising ‘autumn’, ‘snowy’, ‘sunset’, ‘watercolor art’, ‘rainbow’, ‘aurora’, ‘mosaic’, ‘ukiyo-e’ and ‘a sketch with crayon’. For domain adaptation, we introduce an additional set of bespoke conditional prompts, such as ‘graffiti’, ‘retro comic’, ‘chalk drawing’, ‘watercolor painting’, ‘digital art’ and ‘cartoon style’. These prompts are designed to introduce a broader range of visual appearances, enhancing the model’s adaptability to different domains ensuring their generic nature and wide applicability. These prompts do not significantly edit the image structure while introducing global visual effects. Each prompt in the textual library is appended with the template ‘A transformed version of image into $prompt$’ to form a specific input for the image editing diffusion model. Examples of images generated using these prompts are shown in Figure 3.
TABLE II: Top-1 and Top-5 accuracies comparison on ImageNet-1K using ResNet-50, on Tiny ImageNet-200 and CIFAR-100 using PreActResNet-18. FGSM error rates on CIFAR-100 and Tiny-ImageNet-200 datasets are also computed for PreActResNet-18. Compared numbers are taken either from the original papers or from Kang and Kim (2023), following the exact protocols.
Method ImageNet-1K Tiny ImageNet-200 CIFAR-100
Top-1 Acc (%) Top-5 Acc (%) Top-1 Acc (%) Top-5 Acc (%) FGSM Error (%) Top-1 Acc (%) Top-5 Acc (%) FGSM Error (%)
Vanilla-He et al. (2016b) 75.97 92.66 57.23 73.65 42.77 76.33 91.02 23.67
AugMix Hendrycks et al. (2019) 76.75 93.30 55.97 74.68 - 75.31 91.62 43.33
ManifoldMix Verma et al. (2019) 76.85 93.50 58.01 74.12 41.99 79.02 93.37 20.98
Mixup Zhang et al. (2018) 77.03 93.52 56.59 73.02 43.41 76.84 92.42 23.16
CutMix Yun et al. (2019b) 77.08 93.45 56.67 75.52 43.33 76.80 91.91 23.20
Guided-SR Kim et al. (2020a) 77.20 93.66 55.97 74.68 - 80.60 94.00 -
PixMix Hendrycks et al. (2022) 77.40 - - - - 79.70 - -
PuzzleMix Kim et al. (2020a) 77.51 93.76 63.48 75.52 36.52 80.38 94.15 19.62
GuidedMix Kang and Kim (2023) 77.53 93.86 64.63 82.49 - 81.20 94.88 -
Co-Mixup Kim et al. (2020b) 77.63 93.84 64.15 - - 80.15 - -
YOCO Han et al. (2022) 77.88 - - - - - - -
Azizi et al. Azizi et al. (2023) 78.17 - - - - - - -
GMix 78.73 95.47 65.80 83.70 34.47 82.58 95.51 16.83
5 Performance Evaluation
5.1 General Classification
General Classification (GC) is an essential task used to assess the effectiveness of data augmentation approaches. An increase in GC accuracy indicates that the augmentation method fairly introduces meaningful variations in input data to enhance the learning process. In our evaluation, we applied our method to three challenging datasets: CIFAR-100, ImageNet-1K, and Tiny-ImageNet-200. Aligning with the practices of current state-of-the-art (SOTA) methods, we used ResNet50 for the ImageNet-1K, and utilized the PreActResNet18 for the Tiny-ImageNet-200 and CIFAR-100 datasets Huang et al. (2021); Kang and Kim (2023); Kim et al. (2020b); Uddin et al. (2020), and employed ResNet50 for the ImageNet dataset Kim et al. (2020a); Hendrycks et al. (2019); Yun et al. (2019b).
Table II showcases a comparison of Top-1 and Top-5 performances on ImageNet-1K dataset, and compares with the existing SOTA methods. GMix exhibited strong performance, achieving a Top-1 accuracy gain of $2.76\%$ over the Vanilla He et al. (2016b) performer. Compared to GuidedMixup Kang and Kim (2023), which shows the second best results for Top-5 accuracy, our approach attains an absolute performance improvement of $1.6\%$ for Top-5 and $1.2\%$ for Top-1 accuracy. Table II also showcases a comparison of performances on Tiny-ImageNet-200 with existing SOTA methods. Our proposed GMix demonstrates superior performance over existing data augmentation approaches. Over the Vanilla He et al. (2016b), we obtained absolute Top-1 and Top-5 performance gains of $8.57\%$ and $10.05\%$, respectively. We outperform the strong baseline of GuidedMixup Kang and Kim (2023) for Top-5 accuracy by an absolute gain of $1.2\%$.
We observe similar trends on the CIFAR-100 dataset. GMix outperforms the Vanilla He et al. (2016b) with Top-1 and Top-5 absolute accuracy gains of $6.25\%$ and $4.49\%$, respectively. These results across diverse and challenging benchmark datasets underscore the effectiveness of GMix in enhancing learning outcomes. They also indicate its capability to mitigate overfitting and promote better generalization in neural network training.
5.2 Adversarial Robustness
In line with current state-of-the-art methods Kim et al. (2020a); Yun et al. (2019b); Hendrycks et al. (2019); Verma et al. (2019); Uddin et al. (2020); Hong et al. (2021); Kim et al. (2020b); Kang and Kim (2023), we assess the robustness of GMix method against adversarial attacks and input perturbations for general classification. For these evaluations, we employ fast adversarial training Wong et al. (2020) to generate adversarially perturbed input images. The primary objective of these tests is to determine if our augmentation technique offers enhanced resistance to adversarial attacks. We measure performance using the FGSM Wong et al. (2020) error rates as an indicator of robustness against such attacks.
As presented in Table II, our approach achieves an error rate of $16.83\%$ on CIFAR-100, outperforming all other methods in comparison. Similarly, on Tiny ImageNet-200, our method outperforms existing state-of-the-art techniques with an error rate of $34.47\%$. These results highlight that our method maintains strong resilience to adversarial perturbations, surpassing the performance of existing top-performing approaches.
5.3 Fine-Grained Visual Classification
Fine-Grained Visual Classification (FGVC) poses arduous challenge compared to general classification because it requires distinguishing subtle variations between objects within the similar category with broader range. To effectively train a model on such tasks, a generative augmentation approach must carefully preserve these fine-grained details. We assess the performance of GMix on FGVC task using three well-known datasets: CUB Nilsback and Zisserman (2008), Stanford Cars Krause et al. (2013), and Aircraft Maji et al. (2013), employing the ResNet-50 He et al. (2016b) architecture. As demonstrated in Table III, GMix outperforms other competitive methods, showing its strength in enhancing the model’s ability to generalize.
It is notable that GMix achieves the highest accuracy across all the datasets. On the CUB dataset, it achieves an absolute gain of top $2.3\%$ over the strong baseline of Guided-AP Kang and Kim (2023). For the Aircraft dataset, GMix achieves an absolute gain of $1.5\%$. Lastly, on the Stanford Cars dataset, the absolute gain is again more than 1%. This consistent performance gain of GMixcan be attributed to the preservation of subtle but important visual information in the hybrid image constructed in our approach.
TABLE III: Top-1 (%) performance comparison on fine-grained task of ResNet-50.
Method Birds Aircraft Cars
Vanilla  He et al. (2016b) 65.50 80.29 85.52
Simple He et al. (2016b) - 80.13 86.36
Auto Augment Cubuk et al. (2019) - 82.28 88.04
Fast AutoAugment Lim et al. (2019) - 82.20 87.19
DADA Li et al. (2020) - 81.16 87.14
RandAugment Cubuk et al. (2020) - 82.30 87.79
AdaAug Cheung and Yeung (2021) - 82.50 88.49
Mixup Zhang et al. (2018) 71.33 82.38 88.14
CutMix Yun et al. (2019b) 72.58 82.45 89.22
SaliencyMix Uddin et al. (2020) 66.66 83.14 89.04
Guided-SR Kang and Kim (2023) 74.08 83.51 89.23
SnapMix  Huang et al. (2021) 75.53 82.96 90.10
PuzzleMix  Kim et al. (2020a) 74.85 82.66 89.68
Co-Mixup  Kim et al. (2020b) 72.83 83.57 89.53
Guided-AP  Kang and Kim (2023) 77.08 84.32 90.27
GMix 79.42 85.84 91.30
TABLE IV: Top-1 (%) accuracy of GMix on fine-tuning experiments using ImageNet pretrained ResNet-50.
Method Flower102 Aircraft Cars
Vanilla He et al. (2016b) 94.98 81.60 88.08
AA Cubuk et al. (2019) 93.88 83.39 90.82
RA Cubuk et al. (2020) 95.23 82.98 89.28
Fast AA Lim et al. (2019) 96.08 82.56 89.71
AdaAug Cheung and Yeung (2021) 97.19 83.97 91.18
GMix 98.27 85.83 93.47
5.4 Transfer Learning
Transfer learning or fine-tuning, is a prevalent technique for adapting large neural network architectures pre-trained on large-scale datasets to smaller datasets. It requires limited computational resources and helps conducting experiments rapidly. Notably, many augmentation methods Yun et al. (2019b); Zhang et al. (2018); Kang and Kim (2023); Uddin et al. (2020); Hendrycks et al. (2019); Kim et al. (2020a, b); Huang et al. (2021); Verma et al. (2019) have not reported their performance for this popular application. Nevertheless, we evaluate our approach by fine-tuning the baseline model on three datasets Flower102, Aircraft, and Stanford Cars, utilizing ResNet-50 model pre-trained on ImageNet. The results are summarized in Table IV.
On the Flower102 dataset, GMix achieved the highest accuracy of $98.27\%$. A similar trend is evident in the Aircraft dataset, where GMix reached $85.83\%$. In the Stanford Cars dataset, GMix again led with $93.47\%$. GMix significantly surpassed other augmentation techniques, demonstrating its effectiveness in transfer learning scenarios. As compared to the vanilla baselines, our method achieves absolute performance gains of $3.3\%$, $4.2\%$ and $5.4\%$ on the used datasets. It is noteworthy that the accuracy obtained by GMix through fine-tuning (Table IV) is comparable to its performance when trained from scratch on the Aircraft and Cars dataset (Table III). Since fine-tuning requires substantially fewer computational resources than training from scratch, this experiment highlights the practical importance of GMix in deep learning applications.
5.5 Data Scarcity
Data scarcity presents a significant challenge when training deep neural networks. With only a limited number of examples per class, deep networks may struggle to learn meaningful patterns, which can lead to overfitting and a reduced ability to generalize. To mitigate this, augmentation techniques are commonly applied to increase the amount of training data. In this context, we evaluate the performance of ResNet-18 He et al. (2016b) trained with just $10$ images per class from the original Flower102 dataset. As shown in Table V, GMix consistently outperforms other mixup approaches, achieving Top-1 accuracy of $77.23\%$ and Top-5 accuracy of $74.35\%$. Our method is specifically designed to diversify the training dataset. By using custom conditional prompts, GMix artificially expands and enriches the dataset, leading to stronger neural network learning.
TABLE V: Top-1 (%) accuracy of ResNet-18 on the data scarcity task using the Flower102 dataset, where only 10 images per class are present in the dataset.
Method Valid Test
Vanilla He et al. (2016b) 64.48 59.14
Mixup Zhang et al. (2018) 70.55 66.81
CutMix Yun et al. (2019b) 62.68 58.51
SaliencyMix Uddin et al. (2020) 63.23 57.45
Guided-SR Kang and Kim (2023) 72.84 69.31
SnapMix Huang et al. (2021) 65.71 59.79
PuzzleMix Kim et al. (2020a) 71.56 66.71
Co-Mixup Kim et al. (2020b) 68.17 63.20
GuidedMixup Kang and Kim (2023) 74.74 70.44
GMix 77.23 74.35
TABLE VI: Comparison of H-score (%) in the Open Partial Domain Adaptation (OPDA) scenario on the Office-Home dataset, where Clipart (Cl), Art (Ar), Real (Re), and Product (Pr) are different domains. Unified (U) methods are applicable for all scenarios including OPDA, OSDA, and PDA. The SF denotes source data-free, and the CF indicates K-means clustering free model adaption. Our method is shown as +GMix to imply LEAD+GMix. Boldface represents performance gains of GMixover LEAD.
Methods U SF CF Ar2Cl Ar2Pr Ar2Re Cl2Ar Cl2Pr Cl2Re Pr2Ar Pr2Cl Pr2Re Re2Ar Re2Cl Re2Pr Avg
CMU Fu et al. (2020) 56.0 56.9 59.2 67.0 64.3 67.8 54.7 51.1 66.4 68.2 57.9 69.7 61.6
DANCE Saito et al. (2020a) 61.0 60.4 64.9 65.7 58.8 61.8 73.1 61.2 66.6 67.7 62.4 63.7 63.9
DCC Li et al. (2021) 58.0 54.1 58.0 74.6 70.6 77.5 64.3 73.6 74.9 81.0 75.1 80.4 70.2
OVANet Saito and Saenko (2021a) 62.8 75.6 78.6 70.7 68.8 75.0 71.3 58.6 80.5 76.1 64.1 78.9 71.8
GATE Chen et al. (2022b) 63.8 75.9 81.4 74.0 72.1 79.8 74.7 70.3 82.7 79.1 71.5 81.7 75.6
UniOT Chang et al. (2022) 67.3 80.5 86.0 73.5 77.3 84.3 75.5 63.3 86.0 77.8 65.4 81.9 76.6
Source-only - - - 47.3 71.6 81.9 51.5 57.2 69.4 56.0 40.3 76.6 61.4 44.2 73.5 60.9
SHOT-O Liang et al. (2020) 32.9 29.5 39.6 56.8 30.1 41.1 54.9 35.4 42.3 58.5 33.5 33.3 40.7
LEAD Qu et al. (2024) 62.7 78.1 86.4 70.6 76.3 83.4 75.3 60.6 86.2 75.4 60.7 83.7 75.0
+ GMix 63.9 83.0 85.9 76.8 78.9 85.3 78.1 62.2 86.9 79.8 62.9 83.0 77.2
TABLE VII: H-score (%) comparison in the Open-set Domain Adaptation (OSDA) scenario on the Office-Home dataset.
Methods U SF CF Ar2Cl Ar2Pr Ar2Re Cl2Ar Cl2Pr Cl2Re Pr2Ar Pr2Cl Pr2Re Re2Ar Re2Cl Re2Pr Avg
CMU Fu et al. (2020) 55.0 57.0 59.0 59.3 58.2 60.6 59.2 51.3 61.2 61.9 53.5 55.3 57.6
DANCE Saito et al. (2020a) 6.5 9.0 9.9 20.4 10.4 9.2 28.4 12.8 12.6 14.2 7.9 13.2 12.9
DCC Li et al. (2021) 56.1 67.5 66.7 49.6 66.5 64.0 55.8 53.0 70.5 61.6 57.2 71.9 61.7
OVANet Saito and Saenko (2021a) 58.6 66.3 69.9 62.0 65.2 68.6 59.8 53.4 69.3 68.7 59.6 66.7 64.0
Source-only - - - 46.1 63.3 72.9 42.8 54.0 58.7 47.8 36.1 66.2 60.8 45.3 68.2 55.2
SHOT-O Liang et al. (2020) 37.7 41.8 48.4 56.4 39.8 40.9 60.0 41.5 49.7 61.8 41.4 43.6 46.9
LEAD Qu et al. (2024) 60.7 70.8 76.5 61.0 68.6 70.8 65.5 59.8 74.2 64.8 57.7 75.8 67.2
+ GMix 62.8 72.4 76.8 65.1 66.8 71.5 67.3 58.5 74.9 65.7 63.1 79.2 68.7
5.6 Domain Adaptation
When DNNs encounter data from distributions that differ from their training data, a noticeable decline in performance is often observed. To tackle this issue, domain adaptation framework Ganin et al. (2016); Kouw and Loog (2019) has been developed, which utilizes labeled data from source domains to train DNNs for unlabeled target domains using a transductive learning approach. We validate the effectiveness of GMix using LEAD framework Qu et al. (2024) on OfficeHome Venkateswara et al. (2017) dataset. For a fair comparison, we access source-only and augment data with GMix. We experiment with different settings including across partial-set domain adaptation (PDA), open-set domain adaptation (OSDA), and open-partial-set domain adaptation (OPDA). The OPDA results are reported in Table VI, while OSDA and PDA performance evaluations are summarized in Tables VII and VIII, respectively.
Open-Partial Domain Adaptation (OPDA). OPDA Saito and Saenko (2021b); You et al. (2019) falls under the broader scope of universal domain adaptation, where there is no prior knowledge about the label shift, such as common classes or the number of categories in the target domain. Some existing approaches Saito et al. (2020b); Chen et al. (2022c) require simultaneous access to both source and target data which may become impractical due to data protection regulations Voigt and Von dem Bussche (2017). In Table VI, our baseline LEAD Qu et al. (2024) has an average H-score of $75.0\%$, while LEAD + GMix achieves $77.2\%$, representing a significant $2.2\%$ improvement in the average H-score. Particularly, for Ar2Pr, addition of our augmentation strategy improved performance of LEAD Qu et al. (2024) by $4.9\%$, for CL2Ar the improvement is $6.2\%$ and for Re2Ar improvement is $4.4\%$. The results affirm that GMix is an effective data augmentation scheme for open partial domain adaptation scenarios.
Open-set Domain Adaptation (OSDA). In OSDA a significant portion of the target samples comes from new categories that are absent in the source domain. Since these categories do not have specific labels, existing methods often treat them collectively as a single ”unknown” class. In Table VII, LEAD + GMix achieved an average improvement in the H-score of $1.5\%$ compared to the baseline LEAD, underscoring the significance of our proposed augmentation strategy. Particularly, considerable improvements of $4.1\%$ for Cl2Ar, $5.4\%$ for Re2CL and $3.4\%$ for Re2Pr; can be observed by the addition of GMix augmentation in the baseline LEAD Qu et al. (2024) method.
Partial Domain Adaptation (PDA). PDA focuses on adapting a model trained on a source domain to a target domain where only a subset of the classes overlap. We compare the performance of a number of techniques on the Office-Home dataset while adapting LEAD as LEAD + GMixto showcase the gain enabled by our method. Table VIII shows the accuracy for various domain adaptation tasks as well as the average accuracy across all tasks. LEAD + GMix achieved an average improvement in the H-score of $0.5\%$ over the LEAD. Particularly, for Pr2Cl maximum improvement of $9.5\%$, in Re2CL, an improvement of $4.6\%$, in Cl2Pr, an improvement of $1.4\%$ is observed by the addition of GMix augmentation.
TABLE VIII: Accuracy comparison (%) in the Partial Domain Adaptation (PDA) scenario on the Office-Home dataset.
Methods U SF CF Ar2Cl Ar2Pr Ar2Re Cl2Ar Cl2Pr Cl2Re Pr2Ar Pr2Cl Pr2Re Re2Ar Re2Cl Re2Pr Avg
CMU Fu et al. (2020) 50.9 74.2 78.4 62.2 64.1 72.5 63.5 47.9 78.3 72.4 54.7 78.9 66.5
DANCE Saito et al. (2020a) 53.6 73.2 84.9 70.8 67.3 82.6 70.0 50.9 84.8 77.0 55.9 81.8 71.1
DCC Li et al. (2021) 54.2 47.5 57.5 83.8 71.6 86.2 63.7 65.0 75.2 85.5 78.2 82.6 70.9
OVANet Saito and Saenko (2021a) 34.1 54.6 72.1 42.4 47.3 55.9 38.2 26.2 61.7 56.7 35.8 68.9 49.5
GATE Chen et al. (2022b) 55.8 75.9 85.3 73.6 70.2 83.0 72.1 59.5 84.7 79.6 63.9 83.8 74.0
Source-only - - - 45.9 69.2 81.1 55.7 61.2 64.8 60.7 41.1 75.8 70.5 49.9 78.4 62.9
LEAD Qu et al. (2024) 58.2 83.1 87.0 70.5 75.4 83.3 73.7 50.4 83.7 78.3 58.7 83.2 73.8
+ GMix 58.9 76.3 87.7 71.3 76.8 83.4 73.9 59.2 81.9 79.1 63.3 80.6 74.3
TABLE IX: Top-1 (%) accuracy of self-supervised learning methods. Adding GMix yields better performance.
Method Flower102 Stanford Cars Aircraft
MoCo v2 He et al. (2020) 80.31 40.82 51.36
+ GMix 82.36 42.27 53.57
SimSiam Chen and He (2021) 86.93 48.34 40.37
+ GMix 89.42 49.35 42.83
5.7 Self-Supervised Learning
We experimented with self-supervised learning models MoCo v2 He et al. (2020) and SimSiam Chen and He (2021) by incorporating our proposed GMix augmentation method. Table IX showcases the Top-1 accuracy of self-supervised learning methods, comparing the performance of MoCo v2 and SimSiam, both with and without augmentation. On the Flower102, Stanford Cars, and Aircraft datasets, the baseline MoCo v2 achieves accuracies of $80.31\%$, $40.82\%$, and $51.36\%$, respectively. Incorporating GMix delivers absolute gains of $2.05\%$, $1.45\%$ and $2.21\%$ on these datasets. Similarly, SimSiam baseline achieved accuracies of $86.93\%$, $48.34\%$, and $40.37\%$ on the Flower102, Stanford Cars, and Aircraft datasets. Augmenting it with the proposed method gives absolute gains of $2.49\%$, $1.01\%$ and $2.46\%$. These results illustrate the efficacy of GMix, as it consistently boosts the performance across various datasets for self-supervised learning methods. The systematic gains highlight the robustness of our approach, demonstrating its potential to enhance the accuracy of state-of-the-art models in different domains.
6 Analysis and Discussion
In this section, we present an in-depth analysis of the different design choices made in GMix, supporting it by a comprehensive ablation study.
6.1 Grad-CAM Visualization
Gradient-weighted Class Activation Mapping (Grad-CAM) Selvaraju et al. (2017) is used to visualize class-discriminative regions in input images. It works by computing the gradients of the target class with respect to the feature maps in the last convolutional layer, highlighting the important regions contributing to the prediction. This helps in understanding where the model is focusing to make decisions, making it useful for interpretability. Using models trained from scratch, comparison with four existing methods Zhang et al. (2018); Yun et al. (2019b); Kim et al. (2020a); Kang and Kim (2023); Uddin et al. (2020) using Grad-CAM visualization is presented in Figure 4. We observe that the models with existing mixup based methods have less focused region in some images, compared to our proposed GMix. It shows that these methods focus specific object features such as front lights or wheels of the car only. The Grad-CAM saliency map of the model trained using GMix focuses on the full object, implying better generalization. This property can be attributed to the fact that GMix works with a single image and mixes different parts of the same image and its edited versions to augment the data. Our method covers more area of the object while focusing less on the context, whereas the compared approaches are more local within the object and in some cases also have focus outside the object. Thus GMix better retains the structure of the foreground object, resulting in an enlarged focus on the entire object representing better generalization.
Figure 4: Representative Grad-CAM visualizations of saliency maps for Mixup Zhang et al. (2018), CutMix Yun et al. (2019b), PuzzleMix Kim et al. (2020a), GuidedMixup Kang and Kim (2023) methods on random input images of Stanford Cars dataset. We train ResNet50 from scratch for a fair comparison. Saliency maps for GMix are more consistent and better focused on the salient image regions.
6.2 Ablation Studies
To assess the significance of the component of our approach, we performed an ablation study by systematically removing the impact of individual components and recording the performance on ResNet-50 across the Stanford Cars and Flowers-102 datasets, as shown in Table X.
In the table, the baseline He et al. (2016b) using only the original images $I_{i}$ obtains Top-1 and Top-5 accuracies of $85.52\%$ $90.34\%$ on Stanford Cars and $78.83\%$ $94.38\%$ on Flowers102. Introducing fractal interpolation ($F_{v}$) to the input images brings a slight improvement in performance for both datasets. We then eliminate both seamless concatenation ($H_{iju}$) and fractal interpolation ($F_{v}$) by conducting experiments using only the edited images ($\hat{I}_{ij}$) as augmented images for training. This setup aligns more closely with methods in Trabucco et al. (2024); Azizi et al. (2023), which also use diffusion-generated images as augmentations. Our results for this experiment in the table are consistent with the insights of Azizi et al. (2023), suggesting that using generated images directly does not always significantly outperform the vanilla approach. We then perform experiments where edited images are interpolated with fractal images during training. In the Cars dataset, this leads to a slight improvement over the baseline, with Top-1 and Top-5 accuracies of $89.42\%$ and $91.57\%$, respectively. However, the Flowers dataset shows a slight drop in performance, indicating that fractal interpolation is more effective when combined with greater diversity in the training set. Next, we remove the fractal interpolation but retain the seamless concatenation of edited ($\hat{I}_{ij}$) and original ($I_{i}$) images to form hybrid images ($H_{iju}$) for data augmentation. This boosts performance for both datasets. We conjecture that this improvement is due to the combined presence of both generated and original image content in each augmented image, highlighting the importance of the seamless concatenation step in GMix. Finally, when all components, including fractal interpolation, are incorporated, the model achieves its best performance for both datasets. These consistent improvements with each additional component emphasize the effectiveness of the design choices in GMix for data augmentation.
TABLE X: Ablation study using Stanford Cars (Cars) and Flowers102 (Flow) datasets. Top-1 and Top-5 accuracies are reported with different combinations of $I_{i}$: Input image, $\hat{I}_{ij}$: Edited images using prompts $p_{j}$, $H_{iju}$: Hybrid images using random mask $M_{u}$, and $F_{v}$: fractal images used to obtain final interpolated image $A_{ijuv}$.
$I_{i}$
$\hat{I}_{ij}$
$H_{iju}$
$F_{v}$
Cars Top-1 85.52 86.73 87.63 89.42 90.83 91.30
Top-5 90.34 92.38 90.23 91.57 97.56 99.96
Flow Top-1 78.73 78.34 77.38 77.81 79.31 80.27
Top-5 94.38 94.91 93.15 93.24 95.38 95.51
6.3 Augmentation Overhead
We evaluate the computational overhead of various state-of-the-art image augmentation methods, including GMix, in relation to their performance improvements. As defined by Kang et al.Kang and Kim (2023), the augmentation overhead $\mathcal{A_{O}}$ is calculated as:
$\mathcal{A_{O}}=\frac{\mathcal{T}_{aug}-\mathcal{T}_{van}}{\mathcal{T}_{van}}\times 100(\%),$
where $\mathcal{T}_{aug}$ refers to the total time required for image generation and training, and $\mathcal{T}_{van}$ represents the training time of the baseline model without augmentation. Although GMix could leverage parallel processing to speed up image generation, in this experiment, we used sequential processing to ensure a fair comparison. As apparent in Fig. 5, GMix achieves an excellent balance between performance and augmentation overhead, surpassing all other methods in accuracy while maintaining significantly lower overhead compared to Co-Mixup Kim et al. (2020b) and SaliencyMix Uddin et al. (2020). Additionally, GMix can be further optimized by saving the generated images offline, enabling quicker execution of multiple training experiments, which is particularly useful for research and optimization purposes.
Figure 5: Augmentation overhead $\mathcal{A_{O}}$ vs Accuracy (%) plot on CUB-200-2011 dataset with batch size 32.
6.4 Impact of Prompt Selection
Diffusion model outputs are dependent on the input prompts Du et al. (2023). Our approach to designing bespoke conditional prompts focuses on creating prompts that edit images while maintaining their structural integrity, making them applicable across various datasets. As outlined in the main text, we introduce filter-like prompts such as snowy, sunset, and rainbow, demonstrating their effectiveness in training robust classifiers. We also explore some bad prompts that are unsuitable for the image editing in GMix as shown in Figure 6. Descriptive or overly complex prompts tend to produce images that deviate significantly from the original distribution, often leading to unrealistic foregrounds and backgrounds, making them ineffective for training. It further emphasizes the value of our proposed filter-like bespoke conditional prompts, which avoid introducing unwanted alterations to the training data.
6.5 Replacing Concatenation with CutMix and Mixup
In these experiments, we integrate GMix with state-of-the-art image augmentation techniques Yun et al. (2019b); Zhang et al. (2018) at an appropriate component level to assess potential performance improvements. Specifically, we replace our seamless concatenation step with CutMix and Mixup, while keeping the rest of the GMix pipeline unchanged. For the CutMix + GMix approach, we substitute the concatenation step in GMix with the random cropping technique from CutMix. Here, a patch is randomly cropped from the generated image and pasted onto the original image. In the Mixup + GMix variant, we replace the concatenation with pixel-level blending of the original and generated images, as described in Zhang et al. (2018). The results are summarized in Table XI. Using either CutMix Yun et al. (2019b) or Mixup Zhang et al. (2018) yields improvements over the original techniques. While GMix provides consistent gains in these settings, the best performance of $81.30\%$ is achieved when our originally proposed method is used.
Figure 6: Some examples of badly designed prompts: Top row: Original training examples. Bottom row: The corresponding generated images. It demonstrates that using detailed prompts (in blue) leads to poor quality images that are unsuitable for training. The image in the last column from CIFAR100, along with its prompt, produces a black image with no visible output.
TABLE XI: Combining GMix with CutMix and Mixup image augmentation methods by replacing the seamless concatenation step of GMix with the masking techniques CutMix & Mixup.
Method Top-1 (%)
ResNet50505050 He et al. (2016b) 78.73
+ CutMix Yun et al. (2019b) 79.22
+ CutMix Yun et al. (2019b) + GMix 79.58
+ Mixup Zhang et al. (2018) 79.34
+ Mixup Zhang et al. (2018) + GMix 80.20
+ GMix 81.30
6.6 Increasing Smooth Mask Vs Performance
Experiments were also conducted using the Flowers102 dataset to assess how different smooth masking strategies affect the overall performance of the proposed method. The results are reported in Table XII. Employing even a single type of mask - specifically, vertical mask; our approach significantly outperforms the vanilla He et al. (2016b) baseline in terms of accuracy. Introducing both vertical and horizontal masks leads to further improvements. The highest accuracy is achieved when both mask types are combined with random flipping between the positions of input and generated images, enhancing the diversity of the training data. Moreover, the method is compatible with masking techniques from previous studies, such as those in Yun et al. (2019b); Zhang et al. (2018).
TABLE XII: Effects of masking in GMix on Flower102 dataset. All variants yield notably superior results compared to the vanilla ResNet-50. However, the best results are achieved when horizontal (Hor) and vertical (Ver) masks are used with flipping.
Mask Top-1 (%) Top-5 (%)
Vanilla He et al. (2016b) 89.74 94.38
Ver Mask ( ) 94.02 98.42
Hor + Ver Masks ( , ) 94.27 99.03
Hor + Ver + Flipping ( , , , ) 95.37 99.39
7 Conclusion
In this work, a data augmentation technique dubbed as GMix is proposed which is built on pre-trained image editing diffusion models. GMix enhances diversity in the training data while preserving the original semantic content of the input images. It employs editing, smooth concatenation, and fractal blending to produce the final augmented images. GMix has shown performance improvement across multiple tasks including general classification, fine-grained classification, domain adaptation, addressing data scarcity, fine-tuning, self-supervised learning, and adversarial robustness, over several architectures and benchmark datasets including ImageNet-1k, Tiny-ImageNet-200, CIFAR-100, Oxford Flower102, Caltech Birds, Stanford-Cars, FGVC Aircraft, and Office-Home. Across a wide range of experiments, our method has consistently demonstrated improved performance outperforming existing SOTA image augmentation methods. The proposed method has the potential to improve learning of deep models across a wide range of tasks.
Limitations: GMix has two key limitations: (1) It relies on text prompts for image editing, and inappropriate text prompts may result in undesired outputs. To address this, we propose a set of filter-like prompts that can be applied to a wide range of natural images. (2) GMix introduces additional overhead for editing images, but this cost may be considered as a trade off for improved convergence in large-scale learning models. Image editing cost can be effectively handled by generating edited images in advance and employing faster image editing methods.